# Trust-informed Decision-Making Through An Uncertainty-Aware Stacked Neural Networks Framework Case Study in COVID-19 Classification

### v1

@@ VII Conclusion
The results of this study indicate that the MCD setting consistently offers a reliable performance baseline across all pre-trained models, serving as a stable foundation for trust-informed decision-making. While the Ensemble setting enhances precision and recall, as evidenced by improved F1 scores, it can introduce variability that impacts the AUC. The EMCD setting, which combines the strengths of MCD and Ensemble strategies, achieves performance metrics comparable to MCD, effectively balancing robustness and variability while maintaining high reliability across all models. Among the pre-trained models evaluated, the ViT emerges as the top performer across all UQ settings, demonstrating superior capability in generating confident and accurate predictions. The BiT model also shows strong performance, particularly under the MCD and EMCD settings, highlighting its reliability in various scenarios.
Moreover, the selection of an appropriate confidence threshold is crucial. Lower confidence thresholds, which increase the number of predictions flagged for review, may be appropriate for scenarios where the stakes are high, such as critical medical diagnoses. Furthermore, the inclusion of PE as an input to the meta-model significantly enhances its performance, resulting in more accurate identification of trustworthy predictions. This enhancement underscores the importance of PE in refining the trustworthiness of ML models and ensuring more reliable decision-making processes in high-stakes environments. Future research is directed toward exploring the enhancement of base model calibration by leveraging the proposed trust-informed confusion matrix. Such advancements have the potential to reduce the RAR and MRR, while simultaneously improving the CAR, CPR, and TPR.

### v2

@@ VII Conclusion
The evaluation was conducted in two primary stages. First, the performance of U-SNN was compared against the traditional threshold-based method using a series of proposed reliability metrics across multiple confidence thresholds and pre-trained models. The results demonstrated that, at lower confidence thresholds while both approaches produced comparable CR and UR values, the U-SNN framework substantially outperformed the traditional method in reducing confidently incorrect predictions. Specifically, U-SNN consistently exhibited lower FCR and CE values, indicating a stronger ability to prevent confidently incorrect predictions and thereby enhance the reliability of certainty assessments. Moreover, U-SNN achieved lower RR values, suggesting greater efficiency in minimizing unnecessary referrals, which is crucial for reducing operational overhead in real-world scenarios.
A key insight from the comparative analysis was that while the traditional threshold-based method showed competitive performance at higher confidence thresholds (e.g., 0.4), such thresholds may not be practical for high-stakes applications. Lower thresholds, which allow models to be more conservative in assigning certainty, are generally preferable for ensuring reliability. In this context, U-SNN demonstrated a significant advantage, particularly at thresholds below 0.4, where confidently incorrect predictions pose the highest risk.
The second stage of analysis involved an ablation study assessing the impact of including or excluding PE as an input to the meta-model. The findings revealed that including PE consistently enhanced U-SNN’s reliability, particularly by reducing FCR and CE values. Although excluding PE led to slightly higher CR values, this came at the cost of increased confidently incorrect predictions, undermining trust in the system’s outputs.
The proposed U-SNN framework provides a robust and reliable alternative to traditional threshold-based uncertainty quantification methods. By leveraging a meta-model and incorporating prediction entropy, U-SNN significantly reduces the risk of confidently incorrect predictions while optimizing the referral process. These advantages are particularly valuable for high-stakes domains, such as healthcare, where trust and decision reliability are paramount. Moreover, the analysis underscores the importance of selecting lower confidence thresholds to enhance model conservativeness and reliability. Future research is directed toward exploring the enhancement of base model calibration to reduce the UR and RR.